OpenStudy (anonymous):

If you combine 240.0 mL of water at 25.00 °C and 120.0 mL of water at 95.00 °C, what is the final temperature of the mixture? Use 1.00 g/mL as the density of water.

OpenStudy (anonymous):

-q = q -mc(Tf-Ti) = mc(Tf-Ti) The energy lost by one = energy gained by the other--thermochemistry, right?

OpenStudy (anonymous):

You need mass in grams, c, the specific heat capacity of water in J/goC and temperatures in degrees Celsius. Keep in mind that the mixture will reach thermal equilibrium, so the final temperature will be one value.

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

Welcome

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

Okay thanks